1a. What: is the FASTEST radial acceleration that you have ever experienced from moving in circulation motion?

b.describe the circumstances

c. how fast you were moving

d. the radius of the circle that you were moving in

e. the acceleration you experienced in terms of m/s and g.

150+ words needed
Answer #1

Okay, Here is my experience

I have experienced maximum radial acceleration at an amusement park while taking a ride on a roller coaster which had a top speed of around 110 km/h. The roller coaster had a large circular section just after a big vertical drop. The radius of the circular section was about 75 m . It was huge. The speed increased drastically at the end of drop and coaster entered the big loop with increased loop and completed the loop with a lot of adrenaline rushing and gushing in.

Here, Let's find out actually how much radial acceleration did I experience.

a = v2 / r

where v is velocity and r is radius

v = 110 km/h = 30.555 m/s

so,

a = 30.5552 / 75

a = 12.448 m/s2

and in terms of g's

a = 12.448 / 9.8

a = 1.27 g's

this is a good amount of g's . Remember, there is limit of how many g's we can sustain .